# Influence of excitation energy on microscopic quantum pathways for ultrafast charge transfer in van der Waals heterostructures

### v2

@@ Abstract
Efficient charge separation in van der Waals (vdW) heterostructures is crucial for optimizing light-harvesting and detection applications. However, precise control over the microscopic pathways governing ultrafast charge transfer remains an open challenge. These pathways are intrinsically linked to charge transfer states with strongly delocalized wave functions that appear at various momenta in the Brillouin zone. Here, we use time- and angle-resolved photoemission spectroscopy (trARPES) to investigate the possibility of steering carriers through specific charge transfer states in a prototypical WS2-graphene heterostructure. By selectively exciting electron-hole pairs at the K-point and close to the Q-point of WS2 with different pump photon energies, we find that charge separation is faster at higher excitation energies. We attribute this to distinct tunneling mechanisms dictated by the momentum where the initial excitation takes place. Our findings reveal that the microscopic pathways for ultrafast charge transfer in vdW heterostructures are highly sensitive to the excitation energy, an aspect that should be carefully considered when optimizing the performance of optoelectronic devices.
@@ I Introduction
The vast selection of 2D materials available today, such as graphene and monolayer transition metal dichalcogenides, enables the design of novel heterostructures with tailored properties [1, 2, 3]. Efficient absorption of visible light followed by ultrafast charge separation [4, 5, 6, 7, 8, 9, 10] is a key feature of these heterostructures, offering great potential for applications in light harvesting and detection. The driving force behind ultrafast charge separation is the band alignment, where photoexcited electrons and holes rapidly relax to the conduction band minimum and valence band maximum in separate layers. Charge transfer states, formed by hybridization between the vdW-coupled layers, create delocalized wave functions that serve as ultrafast tunnelling channels [11, 12, 13]. These states occur at various momenta in the Brillouin zone but do not contribute equally to charge separation. Based on existing models [11, 13, 14, 15, 16, 17, 18], the efficiency of different tunnelling channels depends on the orbital composition of charge transfer states, their momentum-space distribution, and tunnelling barriers that photoexcited carriers must overcome. Consequently, it should be possible to steer quantum pathways by selectively generating electron-hole pairs at specific momenta.
Here, we explore this concept in a prototypical vdW heterostructure consisting of monolayer WS2 and graphene. We photoexcite the heterostructure with pump photon energies of $\hbar\omega=2.0\,\mathrm{eV}$ and $\hbar\omega=3.1\,\mathrm{eV}$ to excite electron-hole pairs at the K-point and in between the $\Gamma$- and Q-point (also referred to as $\Sigma$- or $\Lambda$-point) of WS2, respectively, and probe the charge transfer dynamics directly in the band structure using time- and angle-resolved photoemission spectroscopy (trARPES) [19]. We find that charge separation is significantly faster for $\hbar\omega=3.1\,\mathrm{eV}$, which we attribute to distinct quantum pathways for hole tunneling at different momenta in the Brillouin zone. Our findings introduce new possibilities for optimizing charge separation in vdW heterostructures, paving the way for more efficient light-harvesting and detection technologies.
@@ II Methods
Sample growth: Pretreated 4H-SiC substrates were H-etched to remove scratches and graphitized in Ar atmosphere. The resulting carbon monolayer with $\left(6\sqrt{3}\times 6\sqrt{3}\right)\mathrm{R}30$° structure was decoupled from the SiC substrate by H-intercalation, yielding quasi-freestanding monolayer graphene [20]. WS2 was then grown by chemical vapour deposition (CVD) from solid WO3 and S precursors [21, 22]. Both graphene and WS2 growth were monitored using Raman spectroscopy, atomic force microscopy and secondary electron microscopy. WS2 was found to grow in the shape of triangular islands with side lengths in the range of $300-700\,\mathrm{nm}$ with twist angles of either 0° or 30° with respect to the graphene layer [13].
@@ III Results
Figure 1a shows an ARPES image of the WS2-graphene heterostructure at negative pump-probe delay before the arrival of the pump pulse. The dotted, dashed, and continuous lines indicate the theoretical band structures of WS2 islands with a twist angle of 0°, WS2 islands with a twist angle of 30° [23] and graphene [24], respectively. The bands are shifted in energy to fit the experimentally observed band alignment and doping level. Orange and blue arrows mark the electronic transitions excited at $\hbar\omega=2.0\,\mathrm{eV}$ and $\hbar\omega=3.1\,\mathrm{eV}$, respectively, in WS2. Here, we will only consider WS2 islands with 0° twist angle. The influence of the twist angle on ultrafast charge separation will be investigated in a separate work.
@@
To gain access to the transient carrier population of different areas in the band structure, we integrate the photocurrent over the areas marked by the colored boxes in Figs. 1b and c. The corresponding pump-probe traces are shown in Fig. 2 together with single-exponential fits yielding the decay times $\tau$. Dashed vertical lines in Fig. 2 indicate the pump-probe delay $t_{\mathrm{max}}$ where the respective pump-probe trace reaches its maximum. The fit results are summarized in Table 1. The charge carrier dynamics inside the Dirac cone (Fig. 2a) are found to be quite similar for $\hbar\omega=2.0\,\mathrm{eV}$ and $\hbar\omega=3.1\,\mathrm{eV}$, respectively, considering that the pump pulses (grey-shaded areas in Fig. 2) were slightly longer for $\hbar\omega=3.1\,\mathrm{eV}$ than for $\hbar\omega=2.0\,\mathrm{eV}$. The population dynamics of the WS2 bands, however, show important differences when comparing the two excitation energies. We find that the gain above the equilibrium position of the WS2 VB (Fig. 2b) reaches its maximum $\approx 250\,\mathrm{fs}$ later for $\hbar\omega=2.0\,\mathrm{eV}$ than for $\hbar\omega=3.1\,\mathrm{eV}$. Further, for $\hbar\omega=2.0\,\mathrm{eV}$ the pump-probe signal of the WS2 CB at K is much bigger than close to Q, indicating that the photoexcited electrons are mainly confined to the K valley (Fig. 3b1). For $\hbar\omega=3.1\,\mathrm{eV}$, however, we find that the population of the WS2 CB at K is smaller than the one close to Q and reaches its maximum $\approx 130\,\mathrm{fs}$ later (Fig. 3b2).
@@
To gain direct access to the timescales for ultrafast charge separation, we extract the related charging-induced band shifts of the WS2 VB and CB as follows [18, 13]. First, we extract energy distribution curves (EDCs) at $k=1.2\,\AA ^{-1}$ from the trAPRES snapshots in Fig. 1 that we fit with an appropriate number of Gaussian peaks to obtain the transient binding energies shown in Fig. 3a. Next, we subtract the transient position of the WS2 VB from the transient position of the WS2 CB yielding the transient band gap $E_{\mathrm{gap}}$ shown in Fig. 3b. Finally, assuming that the band gap changes symmetrically around its center, we add (subtract) $E_{\mathrm{gap}}/2$ to (from) the transient position of the WS2 VB (CB) to obtain the shifts $\Delta E_{\mathrm{charge}}^{\mathrm{WS\textsubscript{2}}}$ in Fig. 3c that we previously attributed to a transient charging of the WS2 layer with excess electrons [18, 13]. The corresponding charging shift of the graphene layer in Fig. 3d is obtained by extracting momentum distribution curves (MDCs) in the energy range between $0.2\,\mathrm{eV}$ and $1.0\,\mathrm{eV}$ that we fit with a Lorentzian. The resulting peak positions are then multiplied with the slope of the Dirac cone to yield the data points in Fig. 3d. All orange and blue curves in Fig. 3 are single-exponential fits, the fit parameters of which are summarized in Table 2. Vertical dashed orange and blue lines indicate the positions $t_{\mathrm{max}}$, where the respective pump-probe signals reach their maximum. Note that the fit for the transient band gap in Fig. 3b is sensitive to the value assumed for the equilibrium gap size (we assumed $2.08\,\mathrm{eV}$) which also affects $\Delta E_{\mathrm{charge}}^{\mathrm{WS\textsubscript{2}}}$ in Fig. 3c. We would like to stress, though, that varying the equilibrium gap size by a reasonable amount has only a minor influence on the data points in Fig. 3c. Considering this, the behavior of the transient WS2 band gap at $k=1.2\,\AA ^{-1}$ is very similar for $\hbar\omega=2.0\,\mathrm{eV}$ and $\hbar\omega=3.1\,\mathrm{eV}$. In contrast to this, both the WS2 and the graphene charging shifts are found to depend on the pump photon energy. We find that both charging shifts reach their maximum $160-340\,\mathrm{fs}$ later for $\hbar\omega=2.0\,\mathrm{eV}$ than for $\hbar\omega=3.1\,\mathrm{eV}$.
@@ IV Discussion
The results obtained for $\hbar\omega=2.0\,\mathrm{eV}$ are in good agreement with previous trARPES results [25, 18, 13] making their interpretation straight forward. The asymmetric population dynamics of the Dirac cone (Fig. 2a1), the gain above the equilibrium position of the WS2 VB (Fig. 2b1) and the charging shifts in Figs. 3c and d provide direct evidence of ultrafast charge separation in the WS2-graphene heterostructure, where hole transfer from WS2 to graphene is much faster than electron transfer. The transient reduction of the band gap in Fig. 3b has been previously attributed to screening of the Coulomb interaction by photoinduced free carriers [26, 27, 28, 29, 30, 31, 19].
@@
The observed differences between excitation at $\hbar\omega=2.0\,\mathrm{eV}$ and $\hbar\omega=3.1\,\mathrm{eV}$, however, deserve further discussion. We start with the observation that, for $\hbar\omega=3.1\,\mathrm{eV}$, the population dynamics in the WS2 CB in Fig. 2c2 are different at K and close to Q. The $3.1\,\mathrm{eV}$ pump pulse directly populates states close to the Q valley (see Fig. 1a). These carriers are then observed to scatter to the K valley within $\approx 130\,\mathrm{fs}$ (see Fig. 2c2) in good agreement with previously reported intervalley scattering times for WS2 and similar TMDCs [32, 33, 34, 35, 36].
@@
Next, we focus on the observation that the gain above the equilibrium position of the WS2 VB (Fig. 2b) and the charging shifts in Fig. 3c and d are found to reach their extrema at earlier time delays for $\hbar\omega=3.1\,\mathrm{eV}$ than for $\hbar\omega=2.0\,\mathrm{eV}$. This indicates that hole transfer is faster for $\hbar\omega=3.1\,\mathrm{eV}$ than for $\hbar\omega=2.0\,\mathrm{eV}$. From [18] it is known that the rates for charge transfer from WS2 to graphene increase with increasing electronic temperature as it becomes easier for the photoexcited carriers to overcome the energy barriers to possible charge transfer states. The transient electronic temperatures for the WS2 layer are difficult to determine due to the small population of the WS2 CB and the limited signal-to-noise ratio. The transient electronic temperature of the carriers inside the Dirac cone of graphene, however, can be extracted by fitting its transient population with a Fermi-Dirac distribution. The resulting electronic temperature is displayed in Fig. 4. We find that, despite differences in excitation energy and pump fluence, the transient electronic temperatures for $\hbar\omega=2.0\,\mathrm{eV}$ and $\hbar\omega=3.1\,\mathrm{eV}$ are very similar. Therefore, we conclude that differences in electronic temperature cannot account for the observation that hole transfer is faster for $\hbar\omega=3.1\,\mathrm{eV}$.
Instead, we propose that the different hole transfer rates for $\hbar\omega=2.0\,\mathrm{eV}$ and $\hbar\omega=3.1\,\mathrm{eV}$ can be attributed to different tunneling pathways via different charge transfer states in the WS2 VB. These charge transfer states have been previously identified using density functional theory calculations [13]. For convenience, we sketch the relevant part of the band structure of the WS2-graphene heterostructure in Fig. 5. Charge transfer states, where the WS2 and graphene bands hybridize, are shown in red. For $\hbar\omega=2.0\,\mathrm{eV}$, photoexcitation occurs at the K valley of WS2. The previously proposed scattering processes resulting in ultrafast charge transfer to the graphene layer are indicated by orange arrows [18, 13]. For $\hbar\omega=2.0\,\mathrm{eV}$, hole transfer is believed to occur via the charge transfer state located at $k>K$ in the WS2 VB. Electrons likely transfer into the graphene layer via the charge transfer state located at $Q<k<K$ in the CB. For $3.1\,\mathrm{eV}$ excitation, electron-hole pairs are initially generated at momenta between $\Gamma$ and Q. We propose that, on their way towards the VB maximum at K, these holes pass by the charge transfer state at $Q<k<K$ in the VB where they efficiently transfer into the graphene layer (blue arrows in Fig. 5). The reason why hole transfer at $Q<k<K$ is faster than at $k>K$ is that the size of the avoided crossing between WS2 VB and graphene Dirac cone is bigger (and therefore hybridization is stronger) at $Q<k<K$ than at $k>K$ [13]. The charge transfer state for electrons inside the WS2 CB is likely the same for $\hbar\omega=2.0\,\mathrm{eV}$ and $\hbar\omega=3.1\,\mathrm{eV}$ consistent with similar population lifetimes in Table 1 as well as similar decay times for the charging shifts in Table 2.

### v3

@@ Abstract
Efficient charge separation in van der Waals (vdW) heterostructures is crucial for optimizing light harvesting and detection applications. However, precise control over the microscopic pathways governing ultrafast charge transfer remains an open challenge. These pathways are intrinsically linked to charge transfer states with strongly delocalized wave functions that appear at various momenta in the Brillouin zone. Here, we use time- and angle-resolved photoemission spectroscopy (trARPES) to investigate the possibility of steering carriers through specific charge transfer states in a prototypical WS2-graphene heterostructure. By selectively exciting electron-hole pairs at the K-point (A-exciton resonance) and close to the Q-point (C-exciton resonance) of WS2 with different pump photon energies, we find that charge separation is faster at higher excitation energies. This behavior is attributed to the fact that absorption at the C-exciton resonance generates electron-hole populations at energies well above the direct band gap. The resulting elevated carrier temperatures open an additional, highly efficient charge-transfer channel for holes in the WS2 valence band, leading to an overall acceleration of interlayer hole transfer for C-exciton excitation. The microscopic insights gained in this work can be leveraged to optimize the performance of vdW heterostructures in optoelectronic devices.
@@ I Introduction
The vast selection of 2D materials available today, such as graphene and monolayer transition metal dichalcogenides, enables the design of novel heterostructures with tailored properties [1, 2, 3]. Such heterostructures commonly feature efficient absorption of visible light followed by ultrafast charge separation [4, 5, 6, 7, 8, 9, 10] with great potential for applications in light harvesting and detection. The driving force behind ultrafast charge separation is the band alignment, where photoexcited electrons and holes rapidly relax to the conduction band minimum and valence band maximum in separate layers. Charge transfer states, formed by hybridization between the vdW-coupled layers, create delocalized wave functions that serve as ultrafast tunneling channels [11, 12, 13]. These states occur at various momenta and energies in the Brillouin zone but do not contribute equally to charge separation. Based on existing models [11, 13, 14, 15, 16, 17, 18], the efficiency of different tunneling channels depends on the strength of the hybridization, the size of the energy barriers that photo-excited carriers must overcome to reach the charge transfer state, as well as the available tunneling phase space. The possibility to steer carriers through specific charge transfer states, e.g. by selectively generating them at particular momenta and with controlled excess energies, remains unexplored.
Here, we investigate this concept in a prototypical vdW heterostructure consisting of monolayer WS2 and graphene. We photoexcite the heterostructure with pump photon energies of $\hbar\omega=2.0\,\mathrm{eV}$ and $\hbar\omega=3.1\,\mathrm{eV}$ to excite electron-hole pairs at the K-point (A-exciton) and between the $\Gamma$- and Q-point (C-exciton) of WS2, respectively, and probe the charge transfer dynamics directly in the band structure using time- and angle-resolved photoemission spectroscopy (trARPES) [19]. We find that charge separation is significantly faster following $\hbar\omega=3.1\,\mathrm{eV}$ excitation, which we attribute to the activation of a second, highly efficient charge transfer channel in the valence band. Our findings introduce new possibilities for optimizing charge separation in vdW heterostructures, paving the way for more efficient light-harvesting and detection technologies.
@@ II Methods
Sample growth: Pretreated 4H-SiC substrates were H-etched to remove scratches and graphitized in Ar atmosphere. The resulting carbon monolayer with $\left(6\sqrt{3}\times 6\sqrt{3}\right)\mathrm{R}30^{\circ}$ structure was decoupled from the SiC substrate by H-intercalation, yielding quasi-freestanding monolayer graphene [20]. WS2 was then grown by chemical vapour deposition (CVD) from solid WO3 and S precursors [21, 22]. Both graphene and WS2 growth were monitored using Raman spectroscopy, atomic force microscopy and secondary electron microscopy. WS2 was found to grow in the shape of triangular islands with side lengths in the range of $300-700\,\mathrm{nm}$ with twist angles of either 0° or 30° with respect to the graphene layer [13].
@@ III Results
Figure 1a shows an ARPES image of the WS2-graphene heterostructure at negative pump-probe delay before the arrival of the pump pulse. The dotted, dashed, and continuous lines indicate the theoretical band structures of WS2 islands with a twist angle of 0°, WS2 islands with a twist angle of 30° [23] and graphene [24], respectively. The bands are shifted in energy to fit the experimentally observed band alignment and doping level. Orange and blue arrows mark the electronic transitions in WS2 excited at $\hbar\omega=2.0\,\mathrm{eV}$ and $\hbar\omega=3.1\,\mathrm{eV}$, respectively [25]. Here, we will only consider WS2 islands with 0° twist angle. The influence of the twist angle on ultrafast charge separation will be investigated in a separate work.
@@
To gain access to the transient carrier population of different parts of the band structure, we integrate the photocurrent over the areas marked by the colored boxes in Figs. 1b and c. The corresponding pump-probe traces are shown in Fig. 2 together with single-exponential fits (see supplemental material) yielding the decay times $\tau$. Dashed vertical lines in Fig. 2 indicate the pump-probe delay $t_{\mathrm{ex}}$ where the respective pump-probe trace reaches its extremum. The fit results are summarized in Table 1. The charge carrier dynamics inside the Dirac cone (Fig. 2a) are found to be quite similar for $\hbar\omega=2.0\,\mathrm{eV}$ and $\hbar\omega=3.1\,\mathrm{eV}$, respectively, considering that the pump-pulse cross-correlation (grey-shaded areas in Fig. 2) was longer for $\hbar\omega=3.1\,\mathrm{eV}$ ($90\,\mathrm{fs}$) than for $\hbar\omega=2.0\,\mathrm{eV}$ ($50\,\mathrm{fs}$). The population dynamics of the WS2 bands, however, show important differences when comparing the two excitation energies. We find that the gain above the equilibrium position of the WS2 VB (Fig. 2b) reaches its maximum $\sim 250\,\mathrm{fs}$ later for $\hbar\omega=2.0\,\mathrm{eV}$ than for $\hbar\omega=3.1\,\mathrm{eV}$. Further, for $\hbar\omega=2.0\,\mathrm{eV}$, the pump-probe signal of the WS2 CB at K is much bigger than close to Q, indicating that the photoexcited electrons are mainly confined to the K valley (Fig. 2c1). For $\hbar\omega=3.1\,\mathrm{eV}$, however, we find that the population of the WS2 CB at K is smaller than the one close to Q and reaches its maximum $\sim 130\,\mathrm{fs}$ later (Fig. 2c2).
To gain direct access to the timescales for ultrafast charge separation, we extract the related charging-induced band shifts of the WS2 VB and CB as follows [18, 13]. First, we extract energy distribution curves (EDCs) at selected momenta from the trAPRES snapshots in Fig. 1 that we fit with an appropriate number of Gaussian peaks (see supplemental material) to obtain the transient binding energies shown in Fig. 3a. Next, we subtract the transient position of the WS2 VB from the transient position of the WS2 CB yielding the transient band gap $E_{\mathrm{gap}}$ shown in Fig. 3b. Finally, assuming that the band gap changes symmetrically around its center, we add (subtract) half of the transient band gap change $\Delta E_{\mathrm{gap}}/2$ to (from) the transient position of the WS2 VB (CB) to obtain the shifts $\Delta E_{\mathrm{charge}}^{\mathrm{WS\textsubscript{2}}}$ in Fig. 3c that we previously attributed to a transient charging of the WS2 layer with excess electrons [18, 13].
@@
The corresponding charging shift of the graphene layer in Fig. 3d is obtained by extracting momentum distribution curves (MDCs) that we fit with a Lorentzian (see supplemental material). The resulting peak positions are averaged and multiplied with the slope of the Dirac cone to yield the data points in Fig. 3d. All orange and blue curves in Fig. 3 are single-exponential fits (see supplemental material), the fit parameters of which are summarized in Table 2. Vertical dashed orange and blue lines indicate the positions $t_{\mathrm{ex}}$, where the respective pump-probe signals reach their extrema. Note that the fit for the transient band gap in Fig. 3b is sensitive to the value assumed for the equilibrium gap size (we assumed $E_{\mathrm{gap}}^{\mathrm{equ}}=2.08\,\mathrm{eV}$) which also affects $\Delta E_{\mathrm{charge}}^{\mathrm{WS\textsubscript{2}}}$ in Fig. 3c. We would like to stress, though, that varying the equilibrium gap size by a reasonable amount has only a minor influence on the data points in Fig. 3c. Considering this, the behavior of the transient WS2 band gap at $k=1.2\,\AA ^{-1}$ is very similar for $\hbar\omega=2.0\,\mathrm{eV}$ and $\hbar\omega=3.1\,\mathrm{eV}$. In contrast to this, both the WS2 and the graphene charging shifts are found to depend on the pump photon energy. We find that both charging shifts reach their extrema $160-340\,\mathrm{fs}$ later for $\hbar\omega=2.0\,\mathrm{eV}$ than for $\hbar\omega=3.1\,\mathrm{eV}$.
@@ IV Discussion
The results obtained for $\hbar\omega=2.0\,\mathrm{eV}$ are in good agreement with previous trARPES results [26, 18, 13] making their interpretation straight forward. The asymmetric population dynamics of the Dirac cone (Fig. 2a1), the gain above the equilibrium position of the WS2 VB (Fig. 2b1) and the charging shifts in Figs. 3c and d provide direct evidence of ultrafast charge separation in the WS2-graphene heterostructure, where hole transfer from WS2 to graphene is much faster than electron transfer. The relevant charge transfer states for the WS2-graphene heterostructure have been previously identified using density functional theory calculations [13]. For convenience, we sketch the relevant part of the band structure in Fig. 4. Charge transfer states, where the WS2 and graphene bands hybridize, are shown in red. For $\hbar\omega=2.0\,\mathrm{eV}$, photoexcitation occurs at the K valley of WS2. The previously proposed scattering processes resulting in ultrafast charge transfer to the graphene layer are indicated by orange arrows [18, 13]. The transient reduction of the band gap in Fig. 3b has been previously attributed to screening of the Coulomb interaction by photoinduced free carriers [27, 28, 29, 30, 31, 32, 19].
@@
The observed differences between excitation at $\hbar\omega=2.0\,\mathrm{eV}$ and $\hbar\omega=3.1\,\mathrm{eV}$, however, deserve further discussion. We start with the observation that, for $\hbar\omega=3.1\,\mathrm{eV}$, the population dynamics in the WS2 CB in Fig. 2c2 are different at K and close to Q. The $3.1\,\mathrm{eV}$ pump pulse directly populates states close to the Q valley (see Fig. 1a) [25]. These carriers are then observed to scatter to the K valley within $\sim 130\,\mathrm{fs}$ (see Fig. 2c2) in good agreement with previously reported intervalley scattering times for WS2 and similar TMDCs [33, 34, 35, 36, 37].
@@
Next, we focus on the observation that the gain above the equilibrium position of the WS2 VB (Fig. 2b) and the charging shifts (Fig. 3c and d) are found to reach their extrema at earlier time delays for $\hbar\omega=3.1\,\mathrm{eV}$ than for $\hbar\omega=2.0\,\mathrm{eV}$. This indicates that hole transfer is faster for $\hbar\omega=3.1\,\mathrm{eV}$ than for $\hbar\omega=2.0\,\mathrm{eV}$. From [18] it is known that the rates for charge transfer from WS2 to graphene increase with increasing electronic temperature as it becomes easier for the photoexcited carriers to overcome the energy barriers to the closest charge transfer states.
The transient electronic temperature for the WS2 layer is difficult to determine due to the limited signal-to-noise ratio. Therefore, we estimate the peak electronic temperatures for electrons and holes in WS2 as described in detail in the supplemental material. We obtain $T_{\mathrm{max}}^{3.1\,\mathrm{eV}}\sim 3700\,\mathrm{K}>T_{\mathrm{max}}^{2.0\,\mathrm{eV}}\sim 1000\,\mathrm{K}$ as the excess energy $\frac{1}{2}(\hbar\omega-E_{\mathrm{gap}})$ of the carriers is considerably larger at $\hbar\omega=3.1\,\mathrm{eV}$. This, indeed, makes it easier for the holes to overcome the energy barrier $\Delta E_{\mathrm{K}}\sim 170\,\mathrm{meV}$ (see Fig. 4). In addition, $T_{\mathrm{max}}^{3.1\,\mathrm{eV}}$ is high enough for the holes to reach the second charge transfer state at $Q<k<K$ in the VB with an energy barrier of $\Delta E_{QK}\sim 410\,\mathrm{meV}$. Hole transfer at $Q<k<K$ is expected to be faster than at $k>K$ because the size of the avoided crossing between WS2 VB and graphene Dirac cone is bigger (and therefore hybridization is stronger) [13]. The excitation and charge transfer channels for $\hbar\omega=3.1\,\mathrm{eV}$ are indicated by blue arrows in Fig. 4.
Note that the photoexcited carrier densities (an estimate is provided in the supplementary material) exceed the Mott threshold [28, 38, 39], where photoexcited excitons rapidly decay into free electron-hole pairs. Therefore, we believe excitonic effects to be irrelevant for interpreting our trARPES data.
Finally, we would like to briefly comment on the possible influence of ultrafast energy transfer between WS2 and graphene. Following photoexcitation of vdW heterostructures, both charge and energy transfer are possible relaxation channels. While the charging shifts in our data provide direct evidence of ultrafast charge separation and thus ultrafast charge transfer, indications for ultrafast energy transfer are less obvious. Note that our data shows none of the indications of Meitner-Auger energy transfer discussed in [40]. A possible contribution of Förster energy transfer [41] is difficult to evaluate from our data. However, we would like to point out that the presence or absence of additional energy transfer does not affect our conclusions regarding ultrafast charge transfer.